OpenStudy (anonymous):

How do you change y=x^2-2x-3 to vertex form? The formula is a(x-h)^2 + k I add 3 to both sides and then I don't know what to do after that.. So now I have y+3=1x^2-2x

OpenStudy (wendy.ivette11714):

@jabez177 @KendrickLamar2014 @Awolflover1

OpenStudy (er.mohd.amir):

x^2-2x+1=(x-1)^2

OpenStudy (anonymous):

So it's 1(x-1)^2-3? @Er.Mohd.AMIR

OpenStudy (er.mohd.amir):

No add -1

OpenStudy (anonymous):

complete the square as above y=1(x^2-2x+1-1)-3

OpenStudy (er.mohd.amir):

1(x-1)^2-4

OpenStudy (anonymous):

Why do I add -1? @Er.Mohd.AMIR

OpenStudy (er.mohd.amir):

for complteing square add 1 and therefore subtract 1 for balence the eq.

OpenStudy (er.mohd.amir):

a2+b2-2ab=(a-b)2 so x2-2.x.1+1-1=(x-1)2-1

OpenStudy (anonymous):

Wait, the equation is x^2-2x-3, plus 3 on each end. y+3=x^2-2x then y+3=1(x-1)^2 but im confused on where the 3 goes.. @Er.Mohd.AMIR

OpenStudy (er.mohd.amir):

y=x^2-2x-3 y=x^2-2x+1-1-3 y=(x^2-2x+1)-4 y=(x-1)^2-4

OpenStudy (er.mohd.amir):

now its clear1?

OpenStudy (anonymous):

Ohhhhhhh. Yes! Thank you! I was so confused on this problem.

OpenStudy (er.mohd.amir):

so its clear now